OpenStudy (anonymous):

plz help just answer will give medal and correct answer i have to get a 60% plz

OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? 4x - y = 1 4x + y = 1 x - 4y = -13 4x - y = -7

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@acxbox22

OpenStudy (kendricklamar2014):

4x - y = -7

OpenStudy (anonymous):

thank you so much! can u help me with a few more if so do u want me to post them here or do u want me to make new question so you can get more medals

OpenStudy (kendricklamar2014):

It's not about the medals, OpenStudy rules say not to post more than one question at a time

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

Solve for x: 2|2x - 2| + 4 = 20. x = 5, x = -3 x = 5, x = -5 x = 3, x = -3 x = -5, x = 3 Solve for x: 2x - 2 > 4x + 6. x < -4 x > -4 x < 4 x > 4